Owl shift summary, Dec 19, 1999

pgouffon@charme.if.usp.br
Sun, 19 Dec 1999 11:13:13 -0200

Shift Summary for owl shift, 16/dec/99

On Shift: Hogan, Philippe

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
Tevatron recovery of yesterday's power outages

========================================================================
Runs ended this shift:

none

Tapes brought to FCC this shift:
(including tapes for online split)

none

========================================================================
Accesses during shift:

open access. One to learn how to turn off CSI low voltage supplies if
temperature goes to high.

========================================================================
Overview of shift:

Boring. Fnalu cluster is still down.

========================================================================
Description of Shift:

0115: Hogan goes home
0400: 1 hour test-drive of the couch. OK but would be better with
yesterday's darkness
0515: Only activity: gas check list

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

ArEth (cent) pressure is zero. New bottles may have to be used very soon.

Flashlights are still unusable with burnt bulbs. Fortunately they were
not needed this time.

--------------------------------------------------------